likeabossssssss:

A right rectangular prism has a length of 20.2 cm, a width of 9.1 cm and a height of 14.3 cm. What is the surface area of the right rectangular prism? Enter your answer in the box.

likeabossssssss:

@KjSaif

KjSaif:

SA=2ab+2bc+2ac SA=2*20.2*9.1+2*9.1*14.3+2*20.2*14.3 SA=364+260.26+577.72 SA=1201.98

likeabossssssss:

thank u

KjSaif:

np

likeabossssssss:

A right rectangular prism has a volume of 2414.68 cm³. The length of the solid is 21.2 cm and the width is 8.5 cm. What is the height of the right rectangular prism? Enter your answer in the box.

KjSaif:

V = l * w * h 2414.68 = 21.2 * 8.5 * h 2414.68 = 180.2h divide both sides by 180.2 13.4 = h h = 13.4 cm

KjSaif:

Answer : 13.4